H2SO4 + NaCl → Na2SO4 + HCl, Given 5.09 g of H2SO4 and 6.33 g of NaCl

1.Calculate the mass (in g) of the excess reagent remaining after the reaction is complete.

2.Calculate the theoretical yield of HCl (in g) from the reaction

3.Determine the percent yield of HCl if 1.91 g of HCl is actually produced from the reaction.
Answer #1

The balanced reaction is:

H2SO2 NaCl Na SO+2 HC

Molar mass of H2SO4 is 98 g.mol-1 , thus moles of it was used (5.09 / 98) mol = 0.0519 mol / /

Molar mass of NaCl is 58.5 g.mol-1 , thus moles of it was used (6.33 / 58.5) mol = 0.1082 mol

1. According to the mole ratio of them (H2SO4 : NaCl = 1:2), moles of NaCl should be twice of that of H2SO4. Thus, moles of NaCl should be (2 x 0.0519) mol = 0.1038 mol. So, NaCl is in excess and after complete reaction it will remain (0.1082 - 0.1038) mol = 0.0044 mol = (0.0044 x 58.5) g = 0.2574 g.

2. H2SO4 is the limiting reagent here. So, yield has to be measured based on amount of it. Since moles of HCl produced is twice the moles of H2SO4. Thus, theoretical moles of HCl should be (2 x 0.0519) mol = 0.1038 mol = (0.1038 x 36.5) g = 3.7887 g.

3. Percentage yield = (amount produced / amount should be produced) x % = (1.91 / 3.7887) x 100% = 50.4 %.